OpenStudy (anonymous):

Solve the equation of log (2x + 4) = log (3x -12) with explanation please.

OpenStudy (amistre64):

log a = log b when a=b

OpenStudy (anonymous):

5x-8 thats the farthest I got

OpenStudy (amistre64):

then you took a wrong turn near albuquerque ....

OpenStudy (amistre64):

2x+4 = 3x-12 does not lose its equality ...

OpenStudy (akashdeepdeb):

@amistre64 But you cannot JUST cancel log!! :D

OpenStudy (amistre64):

log is a one to one function ...

OpenStudy (amistre64):

your not canceling perse; you are simple equating the arguments

OpenStudy (amistre64):

log a = log b, only when a=b

OpenStudy (akashdeepdeb):

Okay. :)

OpenStudy (amistre64):

lets assume that a does not equal b: log a = log b log a - log b = 0 log (a/b) = 0 when a/b = 1 a/b = 1 a = b, which is a constradiction to a not= b

OpenStudy (akashdeepdeb):

Exactly what I was jujst going to type! :D Haha.. :D

OpenStudy (akashdeepdeb):

@Ratedover Did you get it?